hartnn (hartnn):

Efficient (or any) way to integrate this: \(\large \sqrt{a^2cos^2 x+b^2sin^2x}dx\)

hartnn (hartnn):

\(\huge \int \large \sqrt{a^2cos^2 x+b^2sin^2x}dx\)

hartnn (hartnn):

@sauravshakya @siddhantsharan

hartnn (hartnn):

@Zekarias

OpenStudy (aravindg):

did you try competing square?

hartnn (hartnn):

adding and subtracting 2ab sinxcosx ?

OpenStudy (aravindg):

yep thats what came into my ind first not foolproof

hartnn (hartnn):

plz.,try...

OpenStudy (aravindg):

did that work @hartnn

OpenStudy (anonymous):

what if we insert 1-sin^2(x) instead of cos^2(x)

hartnn (hartnn):

\(\sqrt{(acos x+bsinx)^2-2absinxcosx}\) hmm....what next ?

OpenStudy (anonymous):

seems to be an elliptic integral and there is no closed form for elliptic integrals

OpenStudy (anonymous):

maybe

hartnn (hartnn):

yes, i heard same thing earlier....but could not digest that this simple function is not integrable......

hartnn (hartnn):

no closed form...any *open* form ?

OpenStudy (anonymous):

You may try the Weierstrass substitution....

OpenStudy (anonymous):

more simpler\[\int \sqrt{1+\sin^2 x} \ \text{d}x\]but no closed form :)

hartnn (hartnn):

ohh...okk....so no closed form means? do we write it in some new function form ?

OpenStudy (anonymous):

closed form means there is no antiderivative for integrand in terms of elementary functions

hartnn (hartnn):

ok,then.....no use banging our heads,solving this....m closing it.....